An Oil and Gas Company offers a position to Amirul with the starting salary of RM54,000 per year and annual increment of 15%.
(a) Find his salary of fifth years.
(b) Show that his total salary from the first year to the nth years is
RM360,000 (1.15^n - 1)
(c) Find the number of years that he has worked if his total salary is exceed RM3, 654, 670.
1
Expert's answer
2020-12-03T07:36:43-0500
  • According to the question, we have that in n years

"T(n) = 54000+54000(1.15)+54000(1.15)^2+....+54000(1.15)^n"

which is clearly a geometric progression, the formula for the nth term of a geometric progression is

"T(n) = ar^{(n-1)}"

therefore "T(5) = 54000*1.15^4"

"T(5) = RM94446.3375"

  • The formula for the sum of terms in a geometric progression is "S(n) = a\frac{r^n-1}{r-1}" where r = 1.15

therefore S(n) = "54000\frac{1.15^n-1}{1.15-1}" ,hence "S(n)=RM360000(1.15^n-1)"

  • Using the formula derived above, we have that

"3654670=360000(1.15^n-1)" therefore

"1.15^n=\frac{3654670}{360000}+1" , taking the log of both sides of the equation

"n=\frac{\log(11.15)}{\log(1.15)}" therefore n = 17.26 years approximately

for Amirui to exceed RM3654670, then he must work for more than 17.26 years